# Key Words.: Gravitational lensing: strong – Methods: data analysis – Methods: observational – Galaxies: clusters: individual: Perseus1 Introduction

### v1

@@ 7 Conclusions
In this work we have investigated the performance of the Euclid telescope for detecting galaxy-scale gravitational lenses. We have employed 41 experts to carry out a blind visual search for lenses in the Euclid ERO data of the Perseus cluster. The search has yielded three grade A and 13 grade B lens candidates.
We have modelled the 16 candidates, to test if the observed VIS images are consistent with a single background galaxy lensed by a simple, plausible lens. We have obtained convincing models for five of the candidates. Modelling for six of the candidates produced inconclusive results, whilst five of the visual inspection candidates are definitively non-lenses based on our modelling.
We have extrapolated the five candidates with valid lens models to the entire EWS and estimated the number of lenses in it that are discoverable with visual inspection and can be confirmed through modelling of the VIS data. This extrapolation yields lenses in the full EWS. This is broadly consistent with the forecast by C15.
Even though our magnitude cut of effectively removes many of the small Einstein radii objects, the distribution of Einstein radii in our modelled sample is still on the higher side when compared to C15, indicating that we are either unable to identify small Einstein radii systems in the visual inspections, or are unable to model them under our simplifying assumptions. Assuming a step function cutoff in Einstein radius, we inferred the cutoff point to be ; below this, our methodology is likely missing most of the lens candidates. Convolving our inferred cutoff with the Collett (2015) Euclid population, we update the forecast to predict galaxy-scale lenses in the whole EWS (assuming that the same visual inspection discovery and modelling methodology are performed on the entire data set). Downscaling to the visual inspected in this work gives approximately lenses, which is in perfect accordance with the five lens candidates with valid models.
This sample represents the first gravitational lenses reported in this patch of the sky, and some of the first lenses discovered with Euclid data. There is tentative evidence that we are missing small Einstein radius lenses, so future work will be necessary to develop identification techniques targeting low-Einstein radii lens candidates. There are also substantial challenges in how we scale up from this field to the full data set: blind visual inspection of is implausible. Neural networks and citizen science are both likely to be needed to reduce the sample to a manageable size.
@@ Acknowledgements.
This work has made use of the Early Release Observations (ERO) data from the Euclid mission of the European Space Agency (ESA), 2024, https://doi.org/10.57780/esa-qmocze3. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). J. A. A. B., B. C., and F. C. acknowledge support from the Swiss National Science Foundation (SNSF). C. OR thanks the Max Planck Society for support through a Max Planck Lise Meitner Group. C. T. and V. B. acknowledge the INAF grant 2022 LEMON. T. E. C. is funded by a Royal Society University Research Fellowship. This project has received funding from the European Research Council (ERC) under the European Unions Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536).References
@@
- Canameras et al. (2023) Canameras, R., Schuldt, S., Shu, Y., et al. 2023, arXiv:2306.03136
@@
- Cuillandre et al. (2024a) Cuillandre, J.-C., Bertin, E., Bolzonella, M., et al. 2024a, A&A, submitted, arXiv:2405.13496
- Cuillandre et al. (2024b) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2024b, A&A, submitted, arXiv:2405.13501
@@
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, submitted, arXiv:2405.13492
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, submitted, arXiv:2405.13493
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., Achúcarro, A., et al. 2024, A&A, submitted, arXiv:2405.13491
@@
- Gilman et al. (2024) Gilman, D., Birrer, S., Nierenberg, A., & Oh, M. S. H. 2024, MNRAS[arXiv:2403.03253]
- Grespan et al. (2024) Grespan, M., Thuruthipilly, H., Pollo, A., et al. 2024, arXiv:2405.11992
@@
- Pascale et al. (2024) Pascale, M., Frye, B. L., Pierel, J. D. R., et al. 2024, arXiv:2403.18902
@@ Appendix A Expert classifications in visual inspections
We show in this appendix the detailed classifications given by the experts when using the 1-by-1 sequential tool.
Appendix B Other lens candidates
In this appendix, we present the lens candidates discovered during the visual search, but that did not become part of our main sample. These samples are dominated by contaminants, making them more representative of the typical false positives found in lens searches rather than of good lens candidates.
Appendix C Modelling results
The results of modelling all 16 grade A and B candidates are shown in Fig. 20.

### v2

@@ 7 Conclusions
In this work we have investigated the performance of the Euclid telescope for detecting galaxy-scale gravitational lenses. We employed 41 experts to carry out a blind visual search for lenses in the Euclid ERO data of the Perseus cluster. The search yielded 3 grade A and 13 grade B lens candidates.
We modelled the 16 candidates to test whether the observed VIS images are consistent with a single background galaxy lensed by a simple, plausible lens. We have obtained convincing models for five of the candidates. Modelling for six of the candidates produced inconclusive results, whilst five of the visual inspection candidates are definitively non-lenses according to our modelling.
We extrapolated the five candidates with valid lens models to the entire EWS and estimated the number of lenses in it that are discoverable with visual inspection and can be confirmed through modelling of the VIS data. This extrapolation yields lenses in the full EWS. This is broadly consistent with the forecast by C15.
Even though our magnitude cut of effectively removes many of the small-Einstein-radius objects, the distribution of Einstein radii in our modelled sample is still on the higher side when compared to C15, indicating that we are either unable to identify small-Einstein-radius systems in the visual inspections or unable to model them under our simplifying assumptions. Assuming a step function cutoff in Einstein radius, we inferred the cutoff point to be ; below this, our methodology is likely missing most of the lens candidates. Convolving our inferred cutoff with the Collett (2015) Euclid population, we now predict that galaxy-scale lenses will be detected in the whole EWS (assuming that the same visual inspection discovery and modelling methodology are performed on the entire dataset). Down-scaling to the visually inspected in this work gives approximately five lenses, which is in perfect accordance with the five lens candidates with valid models.
This sample represents the first gravitational lenses reported in this patch of the sky, and some of the first lenses discovered with Euclid data. There is tentative evidence that we are missing small-Einstein-radius lenses, so future work will be necessary to develop identification techniques targeting small-Einstein-radius lens candidates. There are also substantial challenges in how we scale up from this field to the full dataset: blind visual inspection of is implausible. Neural networks and citizen science are both likely to be needed to reduce the sample to a manageable size.
@@
Data availability
Supplementary materials showing the lens candidates discovered in the search but not included in the main sample are available on Zenodo, at https://zenodo.org/records/14946028.
@@ Acknowledgements.
This work has made use of the Early Release Observations (ERO) data from the Euclid mission of the European Space Agency (ESA), 2024, https://doi.org/10.57780/esa-qmocze3. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). J. A. A. B., B. C., and F. C. acknowledge support from the Swiss National Science Foundation (SNSF). C. O’R thanks the Max Planck Society for support through a Max Planck Lise Meitner Group. C. T. and V. B. acknowledge the INAF grant 2022 LEMON. T. E. C. is funded by a Royal Society University Research Fellowship. S. H. S. thanks the Max Planck Society for support through the Max Planck Fellowship. This research is supported in part by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LensEra: grant agreement No 945536 and LENSNOVA: grant agreement No 771776).References
- Acevedo Barroso et al. (2025) Acevedo Barroso, J. A., Clément, B., Courbin, F., et al. 2025, arXiv:2503.10610
@@
- Cañameras et al. (2024) Cañameras, R., Schuldt, S., Shu, Y., et al. 2024, A&A, 692, A72
@@
- Cuillandre et al. (2025a) Cuillandre, J.-C., Bertin, E., Bolzonella, M., et al. 2025a, A&A, 697, A6
- Cuillandre et al. (2025b) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2025b, A&A, 697, A11
@@
- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al-Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
@@
- Gilman et al. (2024) Gilman, D., Birrer, S., Nierenberg, A., & Oh, M. S. H. 2024, MNRAS, 533, 1687
- González et al. (2025) González, J., Holloway, P., Collett, T., et al. 2025, arXiv:2501.15679
- Grespan et al. (2024) Grespan, M., Thuruthipilly, H., Pollo, A., et al. 2024, A&A, 688, A34
@@
- Pascale et al. (2025) Pascale, M., Frye, B. L., Pierel, J. D. R., et al. 2025, ApJ, 979, 13
@@ Appendix A Expert classifications in visual inspections
We show in this appendix the detailed classifications given by the experts when using the one-by-one sequential tool.
Appendix B Modelling results